My sister is 10 years older than I am. Our combined age is half the age of our grandfather. Our grandfather is 68 years old. How old am I ?
Answer by addingup(3677)

You and sis = 68/2 = 34
You+10 = sis. Use this value for sis above, like this:
You+you+10 = 34
2you = 24
you = 12
and your sister is 10 years older:
12+10 = 22 is your sister's age
:
substitute you with x and sis with y or whatever sign your teacher likes to use for variables.